# Double Check My Desired Return: Transformer with Target Alignment for Offline Reinforcement Learning

### v1

@@ D.2 EpiCare
Environment. EpiCare (Hargrave, Spaeth, and Grosenick 2024) is a flexible yet clinically motivated benchmark to evaluate offline RL and safety-aware decision making in dynamic treatment regimes. EpiCare frames longitudinal patient care as a finite-horizon POMDP (Kaelbling, Littman, and Cassandra 1998). The hidden state space $S=\left\{s_{r},s_{a},s_{1},\ldots,s_{n_{s}}\right\}$ contains $n_{s}$ disease states plus terminal remission $\left(s_{r}\right)$ and adverse event $\left(s_{a}\right)$ states, the discrete action set $A=\left\{a_{1},\ldots,a_{n_{a}}\right\}$ represents treatments, and the observation space $O=[0,1]^{d_{o}}$ records $d_{o}$ separate symptoms that the clinician sees. Disease progression obeys a transition function $T$ that is multiplicatively modulated by each treatment’s vector $m_{a}$, together with treatment-specific remission probabilities $T\left(s_{r}\mid s_{i},a\right)$ and an observation-triggered adverse-event rule, this yields the transition function:
@@
E Additional Experimental Results
E.1 Visualization Example for Return Alignment
In this section, we provide an intuitive toy demonstration to illustrate the efficacy of return alignment achieved by Doctor. We record the gif on the maze2d-umaze-dense task. The Maze2D dataset contains suboptimal trajectories and is specifically designed to evaluate the stitching ability. Supervised learning-based methods, such as DT, struggle when faced with suboptimal trajectories. As a reference, the normalized return of DT in this environment is -6.8, indicating a complete collapse in performance (Yamagata, Khalil, and Santos-Rodriguez 2023).
@@ E.1 Visualization Example for Return Alignment
E.2 Additional Results for Return Alignment
Figure 8: We evaluate the alignment ability of Doctor on the walker2d-medium-replay-v2 with the top returns of trajectories removed. Doctor achieves much better alignment across a wide range of target returns compared to DT, MTM and QT.
In our additional results as shown in Figure 8, we evaluated the return alignment capabilities of Doctor under conditions where optimal data was removed. Notably, even with the removal of the top 5% of the highest-quality data, traditional RvS methods such as MTM and DT exhibited significant performance degradation, both in terms of extrapolation and interpolation. QT, by incorporating a Q-value regularizer, enhanced its stitching ability and demonstrated robust performance in maximizing the expected return. However, this improvement appeared to come at the cost of reduced interpolation accuracy and struggled to precisely align with the target return. Remarkably, Doctor consistently achieved the best alignment performance, closely adhering to the ideal line and outperforming all other methods in maintaining precise alignment.
E.3 Complex Control Tasks
Adroit Tasks BC CQL IQL TD3+BC DT MTM QT Doctor
[…]
Adroit Sum 37.6 41.7 41.0 5.0 95.1 103.2 124.8 128.5
Table 4: Offline results in the Adroit domain. Doctor achieves competitive or superior results compared to the baselines.
In addition to evaluation in the MuJoCo domain, we further assess Doctor in the challenging Adroit tasks, which consist of three dexterous control tasks, each requiring precise interactions and hand-object coordination. As shown in Table 4, we compare both RvS methods (DT, MTM) and value-based algorithms (CQL, IQL, TD3+BC) as well as the recent combination method QT. In the Pen and Door environment, Doctor outperforms other baselines, demonstrating Doctor’s robustness in a range of tasks with varying features.
E.4 Stitching Ability
Maze2D Tasks BC CQL IQL TD3+BC DT MTM QDT Doctor
[…]
Maze2d Sum 128.7 186.1 138.7 165.5 41.5 76.5 163.1 279.1
Table 5: Offline results in the maze2d domain. Doctor achieves competitive or superior results compared to the baselines.
E.5 Additional Results on EpiCare
Figure 9: Mean returns and number of adverse events per 10k episodes (lower is better) for each offline learning method on the SoC dataset, averaged over eight structurally distinct EpiCare environments (environment seeds 1–8). Error bars denote standard deviation.
Figure 9 reports the mean returns and the number of adverse events (lower is better) for each offline learning method on the SoC dataset, averaged over eight structurally distinct EpiCare environments (environment seeds 1–8). Error bars denote standard deviation. Doctor achieves the lowest adverse events and the best clinical returns, outperforming other baselines and demonstrating that Doctor’s alignment produces substantially safer and more effective treatment policies on the dynamic treatment regime benchmark.
E.6 Error Between the Actual Return and the Target Return
Dataset DT RADT QT Doctor
[…]
$104.3$ $63.8$ $148.6$ $\mathbf{59.2}$
Table 6: Absolute error comparison in the Locomotion tasks. We report the average absolute error over target returns. The results are averaged over three random seeds. Lower summation is better. Doctor achieves comparable performance with RADT in terms of minimizing absolute error.
In this section, we evaluate the absolute error between the actual return and the target return of Doctor. We compare Doctor with other RvS-based baselines. The absolute error is calculated using the following formula:
$E_{\text{diff}}\left(R_{\text{target }},R_{\text{actual }}\right)=\left|R_{\text{target }}-\sum_{t=1}^{T}r_{t}\right|.$ (17)
We report the final absolute error over 3 random seeds in Gym locomotion-v2 tasks as shown in Table 6. The target return is set to 0x-1.0x times the maximum return in the dataset. We put the maximum return of the different dataset in Table 7. According to the table, DT and QT have large absolute error across varying targets. Doctor effectively aligns the actual return with the target return and achieves competitive performance with RADT in terms of minimizing absolute error.
Dataset max $R$ mean $R$ normalized max $R$
hopper-M 3222 1422 99.5
hopper-MR 3192 467 98.6
halfcheetah-M 5309 4770 45.0
halfcheetah-MR 4985 3093 42.4
walker2d-M 4226 2852 92.0
walker2d-MR 4132 682 89.9
Table 7: Scores in Gym locomotion datasets. This table shows maximum return, mean return and normalized return for each dataset.
E.7 Additional Results for Online Finetuning
@@
where $\beta$ is a temperature parameter that controls the sharpness of the distribution. This results in an effective exploration strategy that integrates prior knowledge from the value functions and the desired target returns, as summarized in Alg. 1.
@@
As shown in Table 8, Doctor achieves a performance improvement on some tasks when incorporating additional online interaction data. ODT (Zheng, Zhang, and Grover 2022) is a DT-based variant trained with sequence-level entropy regularization for offline-to-online fine-tuning. MTM-O is an online variant of MTM we implement for test its ability of fintuning. Doctor outperforms ODT and IQL in several tasks, with clear improvements in performance on Hopper-M. This demonstrates the effectiveness of the double-check mechanism for online fine-tuning, allowing the model to explore effectively. The availability of online data can further improve the performance of transformer models.
@@
Tau 0.7 for Locomotion, 0.9 for Maze2d and
0.8 for Adroit and EpiCare
@@
F Model and Training Details
F.1 Model Hyperparameters and Baseline Details
We build our policy and $Q$ value heads as a transformer-based model. The detailed model hyperparameters are in Table 9.
For DT, we train the model using the setting inherit from CORL (Tarasov et al. 2024). We inherit the QT settings from the source code provided by the original paper authors. Model parameters of QT and DT are given in Table 10.
Parameter Value
Number of layers 4
Number of attention heads 4
Embedding dimension 512
Nonlinearity function ReLU
Batch size 256
Context length $K$ 20
Dropout 0.1
Learning rate 3e-4
Table 10: Hyperparameters of QT/DT in experiments.
F.2 Training Details
We provide implementation details of Doctor. The supervised learning model include a bidirectional transformer encoder and a bidirectional transformer decoder. Before inputting the sequence data into the model, each input modality is projected into the embedding space through independent embed-encodings. The output of the decoder is connected to a 2-layer MLP with layer normalization, which is used to reconstruct the trajectory sequence. The Transformer is trained with a randomly sampled series of mask ratios similar to (Wu et al. 2023): mask_ratios = [0.6, 0.7, 0.8, 0.85, 0.9, 0.95, 1.0]. For data sampling, we adopt a two-step sampling method similar to that used in DT (Chen et al. 2021), where we first sample a single trajectory and then uniformly sample sub-trajectories of a certain sequence length.
For offline training, we initialize the AdamW optimizer for the Transformer model and the Adam optimizer for the Q-value head, employing both warmup and decay schedules. The Q-value head consists of two 256-dimensional MLP layers, which connect to the output of the transformer decoder. All hyperparameters are summarized in Table 9.
During the fine-tuning stage, we initialize the replay buffer with the top 5% highest-return trajectories from the offline dataset. Each time we interact with the environment, we fully roll out one episode using the current policy and add it to the replay buffer. We then update the policy and roll out again, following a process similar to (Zheng, Zhang, and Grover 2022).
G Limitations
In this paper, we introduce Doctor to address the return‐alignment problem and demonstrate its effectiveness on standard RL benchmarks such as MuJoCo, Maze2d, Adroit, and EpiCare. Although EpiCare captures many real‐world clinical challenges, such as short treatment horizons, heterogeneous treatment effects, partial observability, and adverse events, it remains a simulated environment. As a next step, we plan to evaluate our method on real‐world clinical datasets.
H Borader Impacts
This study focuses on improving the alignment accuracy of RvS methods in offline reinforcement learning. By addressing their current limitations and achieving a closer match between actual and target returns, it advances the field of offline reinforcement learning. As foundational machine learning research, it poses no negative societal impacts.

### v2

@@ C.2 EpiCare
Environment. EpiCare [5] is a flexible yet clinically motivated benchmark to evaluate offline RL and safety-aware decision making in dynamic treatment regimes. EpiCare frames longitudinal patient care as a finite-horizon POMDP. The hidden state space $S=\left\{s_{r},s_{a},s_{1},\ldots,s_{n_{s}}\right\}$ contains $n_{s}$ disease states plus terminal remission $\left(s_{r}\right)$ and adverse event $\left(s_{a}\right)$ states, the discrete action set $A=\left\{a_{1},\ldots,a_{n_{a}}\right\}$ represents treatments, and the observation space $O=[0,1]^{d_{o}}$ records $d_{o}$ separate symptoms that the clinician sees. Disease progression obeys a transition function $T$ that is multiplicatively modulated by each treatment’s vector $m_{a}$, together with treatment-specific remission probabilities $T\left(s_{r}\mid s_{i},a\right)$ and an observation-triggered adverse-event rule, this yields the transition function:
@@
Appendix D Additional Experimental Results
D.1 Visualization Example for Return Alignment
In this section, we provide an intuitive toy demonstration to illustrate the efficacy of return alignment achieved by Doctor. We record the gif on the maze2d-umaze-dense task. The Maze2D dataset contains suboptimal trajectories and is specifically designed to evaluate the stitching ability. Supervised learning-based methods, such as DT, struggle when faced with suboptimal trajectories. As a reference, the normalized return of DT in this environment is -6.8, indicating a complete collapse in performance [28].
@@ D.1 Visualization Example for Return Alignment
D.2 Additional Results for Online Finetuning
@@
where $\beta$ is a temperature parameter that controls the sharpness of the distribution.
This results in an effective exploration strategy that integrates prior knowledge from the value functions and the desired target returns.
@@
As shown in Table 6, Doctor achieves a performance improvement on some tasks when incorporating additional online interaction data. ODT [39] is a DT-based variant trained with sequence-level entropy regularization for offline-to-online fine-tuning. MTM-O is an online variant of MTM we implement for test its ability of fintuning. Doctor outperforms ODT and IQL in several tasks, with clear improvements in performance on Hopper-M. This demonstrates the effectiveness of the double-check mechanism for online fine-tuning, allowing the model to explore effectively. The availability of online data can further improve the performance of transformer models.
@@
Tau 0.7 (Locomotion), 0.9 (Maze2d),
0.8 (Adroit and EpiCare)
@@
Appendix E Model and Training Details
We build our policy and $Q$ value heads as a transformer-based model. The detailed model hyperparameters are in Table 7. For baselines, we train the model using the setting inherited from CORL [40].
We provide implementation details of Doctor. The supervised learning model include a bidirectional transformer encoder and a bidirectional transformer decoder. Before inputting the sequence data into the model, each input modality is projected into the embedding space through independent embed-encodings. The output of the decoder is connected to a 2-layer MLP with layer normalization, which is used to reconstruct the trajectory sequence. The Transformer is trained with a randomly sampled series of mask ratios similar to [27]: mask_ratios = [0.6, 0.7, 0.8, 0.85, 0.9, 0.95, 1.0]. For data sampling, we adopt a two-step sampling method similar to that used in DT [12], where we first sample a single trajectory and then uniformly sample sub-trajectories of a certain sequence length.
For offline training, we perform 140,000 gradient updates during training, and evaluate the model by rolling out 10 episodes. we set $N=300$ and $\delta$ is small value fluctuating based on the maximum return in the dataset. Specifically, we choose $\delta=5\%\times R_{\text{max}}$.
We initialize the AdamW optimizer for the Transformer model and the Adam optimizer for the Q-value head, employing both warmup and decay schedules. The Q-value head consists of two 256-dimensional MLP layers, which connect to the output of the transformer decoder. All hyperparameters are summarized in Table 7. We report the final results over five random seeds.
During the fine-tuning stage, we initialize the replay buffer with the top 5% highest-return trajectories from the offline dataset. Each time we interact with the environment, we fully roll out one episode using the current policy and add it to the replay buffer. We then update the policy and roll out again, following a process similar to [39].